# Fractal transference principles for subsets of $\mathbb{N}^{d}$ of positive density

### v1

@@ 2.2 Scaling and reduction lemmas in $\mathbb{N}^{d}$
Let $S\subset\mathbb{N}^{d}$ have polynomial density with exponent $\alpha\geq 1$. Then for any $t,L>1$ and all sufficiently large $n$, we have
$\frac{|S\cap Q_{t^{n}}|}{|S\cap Q_{Lt^{n}}|}\asymp L^{-d/\alpha}.$
More precisely, for any $t,L>1$, there exist positive constants $C_{1}^{\prime}$ and $C_{2}^{\prime}$, independent of $L$ and $t$, such that for all sufficiently large $n$:
$C_{1}^{\prime}L^{-d/\alpha}\leq\frac{|S\cap Q_{t^{n}}|}{|S\cap Q_{Lt^{n}}|}\leq C_{2}^{\prime}L^{-d/\alpha}.$
@@
Since $S$ has polynomial density with exponent $\alpha$, there exist constants $C_{1},C_{2}>0$ and $\beta\geq 0$ such that for all sufficiently large $R$:
@@
For the upper bound:
$\frac{|S\cap Q_{t^{n}}|}{|S\cap Q_{Lt^{n}}|}\leq\frac{C_{2}\frac{(t^{n})^{d/\alpha}}{(\log t^{n})^{\beta}}}{C_{1}\frac{(Lt^{n})^{d/\alpha}}{(\log(Lt^{n}))^{\beta}}}=\frac{C_{2}}{C_{1}}L^{-d/\alpha}\cdot\frac{(\log(Lt^{n}))^{\beta}}{(\log t^{n})^{\beta}}.$
For the lower bound:
$\frac{|S\cap Q_{t^{n}}|}{|S\cap Q_{Lt^{n}}|}\geq\frac{C_{1}\frac{(t^{n})^{d/\alpha}}{(\log t^{n})^{\beta}}}{C_{2}\frac{(Lt^{n})^{d/\alpha}}{(\log(Lt^{n}))^{\beta}}}=\frac{C_{1}}{C_{2}}L^{-d/\alpha}\cdot\frac{(\log t^{n})^{\beta}}{(\log(Lt^{n}))^{\beta}}.$
Now observe that:
$\frac{(\log(Lt^{n}))^{\beta}}{(\log t^{n})^{\beta}}=\left(1+\frac{\log L}{n\log t}\right)^{\beta}\to 1\quad\text{as }n\to\infty,$
and similarly:
$\frac{(\log t^{n})^{\beta}}{(\log(Lt^{n}))^{\beta}}=\left(1-\frac{\log L}{n\log t+\log L}\right)^{\beta}\to 1\quad\text{as }n\to\infty.$
Therefore, for sufficiently large $n$, we have:
$\frac{1}{2}\leq\frac{(\log(Lt^{n}))^{\beta}}{(\log t^{n})^{\beta}}\leq 2\quad\text{and}\quad\frac{1}{2}\leq\frac{(\log t^{n})^{\beta}}{(\log(Lt^{n}))^{\beta}}\leq 2.$
Combining these estimates, we obtain:
$\frac{C_{1}}{2C_{2}}L^{-d/\alpha}\leq\frac{|S\cap Q_{t^{n}}|}{|S\cap Q_{Lt^{n}}|}\leq\frac{2C_{2}}{C_{1}}L^{-d/\alpha},$
which proves the desired asymptotic equivalence $\frac{|S\cap Q_{t^{n}}|}{|S\cap Q_{Lt^{n}}|}\asymp L^{-d/\alpha}$. ∎
@@
Step 1: Far-out witnesses.
Fix a sequence $\varepsilon_{k}\downarrow 0$, for instance $\varepsilon_{k}:=1/k$. By the definition of $d^{\ast}(S)=\alpha$, for each $k$ we may choose $R_{k}\in\mathbb{N}$ and $\mathbf{u}_{k}\in\mathbb{N}^{d}$ such that
$\frac{|S\cap Q_{R_{k}}(\mathbf{u}_{k})|}{|Q_{R_{k}}|}>\alpha-\varepsilon_{k}.$
Passing to a subsequence, we may assume $R_{k}\to\infty$ and $R_{k}\geq k^{4}$.
Set $N_{k}:=k$, let $q_{k}:=\lfloor R_{k}/N_{k}\rfloor$, and define $R_{k}^{\prime}:=q_{k}N_{k}$. Then $R_{k}^{\prime}\leq R_{k}$ and $0\leq R_{k}-R_{k}^{\prime}<N_{k}$, hence
@@
Since $R_{k}^{\prime}\geq R_{k}-N_{k}$ and $R_{k}\gg N_{k}$, the error term tends to 0. After passing to a further subsequence (and relabelling), we may assume that
$\frac{|S\cap Q_{R_{k}^{\prime}}(\mathbf{u}_{k})|}{|Q_{R_{k}^{\prime}}|}>\alpha-2\varepsilon_{k}\qquad\text{for all }k.$
Partition $Q_{R_{k}^{\prime}}(\mathbf{u}_{k})$ into the $q_{k}^{d}$ disjoint cubes
$Q_{N_{k}}\bigl(\mathbf{u}_{k}+N_{k}\mathbf{m}\bigr),\qquad\mathbf{m}\in\{0,1,\dots,q_{k}-1\}^{d}.$
Let $M_{k}:=k^{2}$. Call a cube $M_{k}$–interior if its translation vector $\mathbf{u}_{k}+N_{k}\mathbf{m}$ satisfies $\min_{i}\bigl((u_{k})_{i}+N_{k}m_{i}\bigr)\geq M_{k}$. Let $\mathcal{I}_{k}$ be the set of interior indices and $\mathcal{B}_{k}$ its complement. For each coordinate $i$, the number of boundary layers in the $i$th direction is at most $\lceil M_{k}/N_{k}\rceil\leq k+1$, hence
@@
Since $R_{k}\geq k^{4}$ and $N_{k}=k$, we have $q_{k}\geq R_{k}/N_{k}-1\geq k^{3}-1$, so for all large $k$,
$\frac{|\mathcal{B}_{k}|}{q_{k}^{d}}\leq\frac{2d}{k^{2}}.$
For $\mathbf{m}\in\{0,\dots,q_{k}-1\}^{d}$ set
@@
Since $\rho_{\mathbf{m}}\leq 1$,
$\sum_{\mathbf{m}\in\mathcal{I}_{k}}\rho_{\mathbf{m}}\geq q_{k}^{d}(\alpha-2\varepsilon_{k})-|\mathcal{B}_{k}|.$
Dividing by $|\mathcal{I}_{k}|=q_{k}^{d}-|\mathcal{B}_{k}|$ and using $\frac{|\mathcal{B}_{k}|}{q_{k}^{d}}\leq\frac{2d}{k^{2}}$ yields, for all large $k$,
$\frac{1}{|\mathcal{I}_{k}|}\sum_{\mathbf{m}\in\mathcal{I}_{k}}\rho_{\mathbf{m}}\geq\alpha-2\varepsilon_{k}-\frac{2d}{k^{2}}.$
Hence there exists $\mathbf{m}_{k}\in\mathcal{I}_{k}$ such that
$\rho_{\mathbf{m}_{k}}>\alpha-2\varepsilon_{k}-\frac{2d}{k^{2}}.$
Define
$\mathbf{v}_{k}:=\mathbf{u}_{k}+N_{k}\mathbf{m}_{k}.$
Then $\min_{i}(v_{k})_{i}\geq M_{k}=k^{2}$, so $\min_{i}(v_{k})_{i}\to\infty$, and
@@
Step 2: Constructing $S^{\prime}$.
We now construct $S^{\prime}\subset S$ supported on a rapidly separated family of cubes. Fix a sequence $\varepsilon_{j}\downarrow 0$ (relabel if necessary). We construct pairwise disjoint cubes $B_{j}=Q_{T_{j}}(\mathbf{w}_{j})$ inductively and set
@@
Assume $B_{1},\dots,B_{j-1}$ have been chosen and put
$H_{j}:=\max_{1\leq i\leq j-1}\bigl(\|\mathbf{w}_{i}\|_{\infty}+T_{i}\bigr)\qquad(\text{with }H_{1}:=1).$
Choose $k_{j}$ so large that
$N_{k_{j}}\geq 2H_{j},\qquad\min_{1\leq i\leq d}(v_{k_{j}})_{i}\geq 10^{j}H_{j},\qquad\frac{|S\cap Q_{N_{k_{j}}}(\mathbf{v}_{k_{j}})|}{|Q_{N_{k_{j}}}|}>\alpha-\varepsilon_{j},$
and additionally
$\frac{2^{d}d\,H_{j}}{N_{k_{j}}}\leq\frac{\varepsilon_{j}}{100}.$
Write $N_{k_{j}}=q_{j}H_{j}+r_{j}$ with $0\leq r_{j}<H_{j}$ and set $N^{\prime}_{j}:=N_{k_{j}}-r_{j}$. Then $N^{\prime}_{j}$ is a multiple of $H_{j}$ and
$|Q_{N_{k_{j}}}|-|Q_{N^{\prime}_{j}}|=N_{k_{j}}^{d}-(N^{\prime}_{j})^{d}\leq d\,r_{j}\,N_{k_{j}}^{d-1}\leq d\,H_{j}\,N_{k_{j}}^{d-1}.$
Hence
$|S\cap Q_{N^{\prime}_{j}}(\mathbf{v}_{k_{j}})|\geq|S\cap Q_{N_{k_{j}}}(\mathbf{v}_{k_{j}})|-\bigl(|Q_{N_{k_{j}}}|-|Q_{N^{\prime}_{j}}|\bigr)>(\alpha-\varepsilon_{j})N_{k_{j}}^{d}-dH_{j}N_{k_{j}}^{d-1}.$
Dividing by $|Q_{N^{\prime}_{j}}|=(N^{\prime}_{j})^{d}$ and using $(N^{\prime}_{j})^{d}\geq(N_{k_{j}}-H_{j})^{d}\geq(N_{k_{j}}/2)^{d}$ (since $N_{k_{j}}\geq 2H_{j})$, we obtain
$\frac{|S\cap Q_{N^{\prime}_{j}}(\mathbf{v}_{k_{j}})|}{|Q_{N^{\prime}_{j}}|}>(\alpha-\varepsilon_{j})-\frac{2^{d}d\,H_{j}}{N_{k_{j}}}\geq\alpha-\varepsilon_{j}-\frac{\varepsilon_{j}}{100}.$
Now partition $Q_{N^{\prime}_{j}}(\mathbf{v}_{k_{j}})$ into disjoint cubes of side length $H_{j}$. By averaging, there exists a cube $Q_{H_{j}}(\mathbf{w}_{j})\subset Q_{N^{\prime}_{j}}(\mathbf{v}_{k_{j}})$ such that
$\frac{|S\cap Q_{H_{j}}(\mathbf{w}_{j})|}{|Q_{H_{j}}|}\geq\frac{|S\cap Q_{N^{\prime}_{j}}(\mathbf{v}_{k_{j}})|}{|Q_{N^{\prime}_{j}}|}>\alpha-\varepsilon_{j}-\frac{\varepsilon_{j}}{100}.$
Set $T_{j}:=H_{j}$ and $B_{j}:=Q_{T_{j}}(\mathbf{w}_{j})$.
Moreover, since $B_{j}\subset Q_{N^{\prime}_{j}}(\mathbf{v}_{k_{j}})$, we have $\min_{i}(w_{j})_{i}\geq\min_{i}(v_{k_{j}})_{i}\geq 10^{j}T_{j}$. On the other hand, the union $\bigcup_{i=1}^{j-1}B_{i}$ is contained in $Q_{H_{j}}=Q_{T_{j}}$ by the definition of $H_{j}$. Therefore $B_{j}\cap\bigcup_{i=1}^{j-1}B_{i}=\emptyset$, so the cubes $(B_{j})$ are pairwise disjoint.
Step 3: Verifying the conclusions.
For each $j$ we have $S^{\prime}\cap B_{j}=S\cap B_{j}$, hence
$d^{\ast}(S^{\prime})\geq\limsup_{j\to\infty}\frac{|S^{\prime}\cap B_{j}|}{|B_{j}|}=\limsup_{j\to\infty}\frac{|S\cap B_{j}|}{|B_{j}|}\geq\alpha.$
Since $S^{\prime}\subset S$, also $d^{\ast}(S^{\prime})\leq d^{\ast}(S)=\alpha$, hence $d^{\ast}(S^{\prime})=\alpha$.
To prove $\overline{d}(S^{\prime})=0$, let $N$ be large and let $j(N)$ be the largest index such that $B_{j(N)}\cap Q_{N}\neq\emptyset$ (if no such index exists, then $S^{\prime}\cap Q_{N}=\emptyset$). Then $B_{m}\cap Q_{N}=\emptyset$ for all $m>j(N)$, and therefore
$|S^{\prime}\cap Q_{N}|\leq\sum_{m=1}^{j(N)}|B_{m}|\leq j(N)\,T_{j(N)}^{d},$
since $T_{m}\leq T_{j(N)}$ for $m\leq j(N)$ by construction. Moreover, $B_{j(N)}\cap Q_{N}\neq\emptyset$ implies $N\geq\min_{i}(w_{j(N)})_{i}$, while $\min_{i}(w_{j(N)})_{i}\geq 10^{j(N)}T_{j(N)}$. Hence
$\frac{|S^{\prime}\cap Q_{N}|}{|Q_{N}|}\leq\frac{j(N)\,T_{j(N)}^{d}}{N^{d}}\leq\frac{j(N)\,T_{j(N)}^{d}}{\min_{i}(w_{j(N)})_{i}^{d}}\leq j(N)\,10^{-d\,j(N)}\xrightarrow[N\to\infty]{}0,$
since $j(N)\to\infty$ as $N\to\infty$. Thus $\overline{d}(S^{\prime})=0$.
Finally, define the sequences $(N_{k})$ and $(\mathbf{v}_{k})$ by
$(N_{k},\mathbf{v}_{k}):=(T_{k},\mathbf{w}_{k}).$
Then $\min_{i}(v_{k})_{i}=\min_{i}(w_{k})_{i}\to\infty$ by construction, and
$\frac{|S^{\prime}\cap Q_{N_{k}}(\mathbf{v}_{k})|}{|Q_{N_{k}}|}=\frac{|S\cap B_{k}|}{|B_{k}|}\longrightarrow\alpha,$
because $|S\cap B_{k}|/|B_{k}|>\alpha-\varepsilon_{k}-\varepsilon_{k}/100$ and $\varepsilon_{k}\to 0$. This completes the proof. ∎
@@
Write $|Q_{N}|=N^{d}$ and set $M:=\lfloor N/K\rfloor$. For $\mathbf{a}=(a_{1},\dots,a_{d})\in Q_{N}$ let $m:=\min_{j}a_{j}$. The condition $\mathbf{a}\in\mathcal{C}_{K}$ is equivalent to $\max_{j}a_{j}\leq Km$. Hence, for each $m\in\{1,\dots,N\}$, all coordinates satisfy
$a_{j}\in[m,U_{m}]\cap\mathbb{N},\qquad U_{m}:=\min\{Km,N\}.$
Let $L_{m}:=|([m,U_{m}]\cap\mathbb{N})|=U_{m}-m+1$. The number of $\mathbf{a}\in Q_{N}$ with $\min_{j}a_{j}=m$ and $\max_{j}a_{j}\leq Km$ equals
$L_{m}^{d}-(L_{m}-1)^{d},$
since $L_{m}^{d}$ counts all dd–tuples in $m,U_{m}]^{d}$ and $(L_{m}-1)^{d}$ counts those with all coordinates at least $m+1$. Therefore,
$|\mathcal{C}_{K}\cap Q_{N}|=\sum_{m=1}^{N}\bigl(L_{m}^{d}-(L_{m}-1)^{d}\bigr).$ (6)
If $1\leq m\leq M$, then $Km\leq N$ so $U_{m}=Km$ and thus $L_{m}=(K-1)m+1$. If $M+1\leq m\leq N$, then $U_{m}=N$ and thus $L_{m}=N-m+1$. Splitting the sum in (6) gives
$|\mathcal{C}_{K}\cap Q_{N}|$ $=\sum_{m=1}^{M}\Bigl(((K-1)m+1)^{d}-((K-1)m)^{d}\Bigr)+\sum_{m=M+1}^{N}\Bigl((N-m+1)^{d}-(N-m)^{d}\Bigr).$ (7)
The second sum telescopes: with t=Nm+1t=N-m+1 it becomes
$\sum_{t=1}^{N-M}\bigl(t^{d}-(t-1)^{d}\bigr)=(N-M)^{d}.$
Hence,
$|\mathcal{C}_{K}\cap Q_{N}|=\sum_{m=1}^{M}\Bigl(((K-1)m+1)^{d}-((K-1)m)^{d}\Bigr)+(N-M)^{d}.$ (8)
For $x\geq 0$, the binomial expansion yields
$(x+1)^{d}-x^{d}=dx^{d-1}+O_{d}(x^{d-2})\qquad(d\geq 2),$
and for $d=1$ the statement is trivial. Applying this with x=(K1)mx=(K-1)m and summing over $m\leq M$ gives
$\sum_{m=1}^{M}\Bigl(((K-1)m+1)^{d}-((K-1)m)^{d}\Bigr)=d(K-1)^{d-1}\sum_{m=1}^{M}m^{d-1}+O_{d,K}\Bigl(\sum_{m=1}^{M}m^{d-2}\Bigr).$
Using the standard power-sum estimates
$\sum_{m=1}^{M}m^{d-1}=\frac{M^{d}}{d}+O_{d}(M^{d-1}),\qquad\sum_{m=1}^{M}m^{d-2}=O_{d}(M^{d-1}),$
we obtain
$\sum_{m=1}^{M}\Bigl(((K-1)m+1)^{d}-((K-1)m)^{d}\Bigr)=(K-1)^{d-1}M^{d}+O_{d,K}(M^{d-1}).$ (9)
Combining (8) and (9) yields
$|\mathcal{C}_{K}\cap Q_{N}|=(K-1)^{d-1}M^{d}+(N-M)^{d}+O_{d,K}(N^{d-1}).$
Dividing by $N^{d}$ and noting that $M/N\to 1/K$ as $N\to\infty$, we get
$\lim_{N\to\infty}\frac{|\mathcal{C}_{K}\cap Q_{N}|}{N^{d}}=(K-1)^{d-1}\Bigl(\frac{1}{K}\Bigr)^{d}+\Bigl(1-\frac{1}{K}\Bigr)^{d}.$
Finally,
$(K-1)^{d-1}\Bigl(\frac{1}{K}\Bigr)^{d}=\Bigl(1-\frac{1}{K}\Bigr)^{d-1}\cdot\frac{1}{K},$
so the limit equals
$\Bigl(1-\frac{1}{K}\Bigr)^{d-1}\Bigl(\frac{1}{K}+1-\frac{1}{K}\Bigr)=\Bigl(1-\frac{1}{K}\Bigr)^{d-1},$
as claimed. ∎
@@
Set $\delta:=\overline{d}(A)>0$. By Lemma 2.12 we have
$\lim_{N\to\infty}\frac{|\mathcal{C}_{K}\cap Q_{N}|}{|Q_{N}|}=\Bigl(1-\frac{1}{K}\Bigr)^{d-1},$
and hence
Choosing $K$ sufficiently large, we may ensure that
$1-\Bigl(1-\frac{1}{K}\Bigr)^{d-1}<\frac{\delta}{4}.$ (10)
By the definition of the limit, there exists $N_{1}$ such that for all $N\geq N_{1}$,
$\frac{|Q_{N}\setminus\mathcal{C}_{K}|}{|Q_{N}|}<\frac{\delta}{2}.$ (11)
By the definition of $\limsup$, there exists a sequence $N_{i}\to\infty$ such that
$\frac{|A\cap Q_{N_{i}}|}{|Q_{N_{i}}|}\longrightarrow\delta.$
In particular, for $i$ sufficiently large we have
$\frac{|A\cap Q_{N_{i}}|}{|Q_{N_{i}}|}>\frac{3\delta}{4},\qquad N_{i}\geq N_{1}.$ (12)
For such $i$, using the decomposition
$A\cap Q_{N_{i}}=\bigl(A\cap\mathcal{C}_{K}\cap Q_{N_{i}}\bigr)\;\cup\;\bigl(A\cap(Q_{N_{i}}\setminus\mathcal{C}_{K})\bigr),$
we obtain
$|A\cap\mathcal{C}_{K}\cap Q_{N_{i}}|\geq|A\cap Q_{N_{i}}|-|Q_{N_{i}}\setminus\mathcal{C}_{K}|.$
Dividing by $|Q_{N_{i}}|$ and combining (11) and (12), we conclude that
$\frac{|A\cap\mathcal{C}_{K}\cap Q_{N_{i}}|}{|Q_{N_{i}}|}\geq\frac{3\delta}{4}-\frac{\delta}{2}=\frac{\delta}{4}.$
Taking the $\limsup$ over $i$ yields
$\overline{d}\bigl(A\cap\mathcal{C}_{K}\bigr)\geq\frac{\delta}{4}>0,$
as claimed. ∎

### v2

@@ 2.2 Scaling and reduction lemmas in $\mathbb{N}^{d}$
If $S\subset\mathbb{N}^{d}$ has polynomial density of exponent $\alpha\geq 1$, then for all $t,L>1$ and all large $n$,
$\frac{|S\cap Q_{t^{n}}|}{|S\cap Q_{Lt^{n}}|}\asymp L^{-d/\alpha},$
with implied constants independent of $t$ and $L$.
@@
By polynomial density, for all large $R$,
@@
Apply this with $R=t^{n}$ and $R=Lt^{n}$ to get
$\frac{C_{1}}{C_{2}}\,L^{-d/\alpha}\Bigl(\frac{\log t^{n}}{\log(Lt^{n})}\Bigr)^{\beta}\leq\frac{|S\cap Q_{t^{n}}|}{|S\cap Q_{Lt^{n}}|}\leq\frac{C_{2}}{C_{1}}\,L^{-d/\alpha}\Bigl(\frac{\log(Lt^{n})}{\log t^{n}}\Bigr)^{\beta}.$
Since $\log(Lt^{n})=n\log t+\log L$, we have
$\Bigl(\frac{\log(Lt^{n})}{\log t^{n}}\Bigr)^{\beta}=\Bigl(1+\frac{\log L}{n\log t}\Bigr)^{\beta}\to 1\quad(n\to\infty),$
and similarly for its reciprocal. Hence for all sufficiently large $n$ these log-factors are bounded above and below by absolute constants (depending only on $\beta$), yielding
$C_{1}^{\prime}L^{-d/\alpha}\leq\frac{|S\cap Q_{t^{n}}|}{|S\cap Q_{Lt^{n}}|}\leq C_{2}^{\prime}L^{-d/\alpha},$
with $C_{1}^{\prime},C_{2}^{\prime}$ independent of $t,L$. ∎
@@
Fix $\varepsilon_{k}:=1/k$ and set $N_{k}:=k$, $M_{k}:=k^{2}$. By $d^{\ast}(S)=\alpha$, choose $R_{k}\in\mathbb{N}$ and $\mathbf{u}_{k}\in\mathbb{N}^{d}$ such that
$\frac{|S\cap Q_{R_{k}}(\mathbf{u}_{k})|}{|Q_{R_{k}}|}>\alpha-\varepsilon_{k},$
and pass to a subsequence so that $R_{k}\geq k^{4}$.
Step 1: far–out witnesses. Let $q_{k}:=\lfloor R_{k}/N_{k}\rfloor$ and $R_{k}^{\prime}:=q_{k}N_{k}$, so $0\leq R_{k}-R_{k}^{\prime}<N_{k}$ and
@@
Since $R_{k}^{\prime}\geq R_{k}-N_{k}$ and $R_{k}\geq k^{4}$, the last term is $\leq\varepsilon_{k}$ for all large $k$. Thus, after discarding finitely many $k$,
$\frac{|S\cap Q_{R_{k}^{\prime}}(\mathbf{u}_{k})|}{|Q_{R_{k}^{\prime}}|}>\alpha-2\varepsilon_{k}.$
Partition $Q_{R_{k}^{\prime}}(\mathbf{u}_{k})$ into the $q_{k}^{d}$ disjoint cubes $Q_{N_{k}}(\mathbf{u}_{k}+N_{k}\mathbf{m})$, $\mathbf{m}\in\{0,\dots,q_{k}-1\}^{d}$. Let $\mathcal{I}_{k}$ be the set of $M_{k}$–interior indices, i.e. $\min_{i}((u_{k})_{i}+N_{k}m_{i})\geq M_{k}$, and let $\mathcal{B}_{k}:=\mathcal{I}_{k}^{c}$. For each coordinate $i$, there are at most $\lceil M_{k}/N_{k}\rceil\leq k+1$ boundary layers, hence
@@
Since $q_{k}\geq R_{k}/N_{k}-1\geq k^{3}-1$, we have $|\mathcal{B}_{k}|/q_{k}^{d}\leq 2d/k^{2}$ for all large $k$.
For $\mathbf{m}$ define
@@
Using $\sum_{\mathbf{m}\in\mathcal{B}_{k}}\rho_{\mathbf{m}}\leq|\mathcal{B}_{k}|$ and $|\mathcal{I}_{k}|=q_{k}^{d}-|\mathcal{B}_{k}|$, we get
$\frac{1}{|\mathcal{I}_{k}|}\sum_{\mathbf{m}\in\mathcal{I}_{k}}\rho_{\mathbf{m}}\geq\frac{q_{k}^{d}(\alpha-2\varepsilon_{k})-|\mathcal{B}_{k}|}{q_{k}^{d}-|\mathcal{B}_{k}|}\geq\alpha-2\varepsilon_{k}-\frac{|\mathcal{B}_{k}|}{q_{k}^{d}-|\mathcal{B}_{k}|}\geq\alpha-2\varepsilon_{k}-\frac{4d}{k^{2}}$
for all large $k$. Hence there exists $\mathbf{m}_{k}\in\mathcal{I}_{k}$ with
$\rho_{\mathbf{m}_{k}}\geq\alpha-2\varepsilon_{k}-\frac{4d}{k^{2}}.$
Set $\mathbf{v}_{k}:=\mathbf{u}_{k}+N_{k}\mathbf{m}_{k}$. Then $\min_{i}(v_{k})_{i}\geq M_{k}\to\infty$ and
@@
Step 2: define $S^{\prime}$. Choose an increasing subsequence $k_{j}$ such that the cubes
$B_{j}:=Q_{N_{k_{j}}}(\mathbf{v}_{k_{j}})$
are pairwise disjoint and satisfy $\min_{i}(\mathbf{v}_{k_{j}})_{i}\geq 10^{j}N_{k_{j}}$ (this is possible since $\min_{i}(\mathbf{v}_{k})_{i}\to\infty$). Define
@@
Step 3: verify densities. Since $S^{\prime}\cap B_{j}=S\cap B_{j}$,
$d^{\ast}(S^{\prime})\geq\limsup_{j\to\infty}\frac{|S^{\prime}\cap B_{j}|}{|B_{j}|}=\limsup_{j\to\infty}\frac{|S\cap B_{j}|}{|B_{j}|}=\alpha,$
and $S^{\prime}\subset S$ gives $d^{\ast}(S^{\prime})\leq d^{\ast}(S)=\alpha$, hence $d^{\ast}(S^{\prime})=\alpha$.
For $\overline{d}(S^{\prime})$, let $N$ be large and let $j=j(N)$ be maximal with $B_{j}\cap Q_{N}\neq\emptyset$. Then $S^{\prime}\cap Q_{N}$ meets only $B_{1},\dots,B_{j}$, so
$|S^{\prime}\cap Q_{N}|\leq\sum_{m\leq j}|B_{m}|\leq j\,N_{k_{j}}^{d}.$
Moreover $B_{j}\cap Q_{N}\neq\emptyset$ implies $N\geq\min_{i}(\mathbf{v}_{k_{j}})_{i}-N_{k_{j}}\geq\frac{1}{2}\min_{i}(\mathbf{v}_{k_{j}})_{i}$ for all large $j$, hence
$\frac{|S^{\prime}\cap Q_{N}|}{|Q_{N}|}\leq\frac{j\,N_{k_{j}}^{d}}{N^{d}}\leq\frac{j\,N_{k_{j}}^{d}}{(\frac{1}{2}\min_{i}(\mathbf{v}_{k_{j}})_{i})^{d}}\leq 2^{d}j\Bigl(\frac{N_{k_{j}}}{\min_{i}(\mathbf{v}_{k_{j}})_{i}}\Bigr)^{d}\leq 2^{d}j\,10^{-dj}\xrightarrow[N\to\infty]{}0.$
Thus $\overline{d}(S^{\prime})=0$, and the cubes $B_{j}$ witness $|S^{\prime}\cap B_{j}|/|B_{j}|\to\alpha$ with $\min_{i}(\mathbf{v}_{k_{j}})_{i}\to\infty$. ∎
@@
Write $|Q_{N}|=N^{d}$ and put $M:=\lfloor N/K\rfloor$. For $\mathbf{a}\in Q_{N}$ let $m:=\min_{j}a_{j}$. The condition $\mathbf{a}\in\mathcal{C}_{K}$ is equivalent to $\max_{j}a_{j}\leq Km$, i.e. $\mathbf{a}\in[m,U_{m}]^{d}$ with $U_{m}:=\min\{Km,N\}$. Let $L_{m}:=U_{m}-m+1$. Then the number of $\mathbf{a}\in\mathcal{C}_{K}\cap Q_{N}$ with $\min_{j}a_{j}=m$ equals $L_{m}^{d}-(L_{m}-1)^{d}$, hence
$|\mathcal{C}_{K}\cap Q_{N}|=\sum_{m=1}^{N}\bigl(L_{m}^{d}-(L_{m}-1)^{d}\bigr).$
If $m\leq M$ then $U_{m}=Km$ and $L_{m}=(K-1)m+1$; if $m>M$ then $U_{m}=N$ and $L_{m}=N-m+1$. Splitting the sum and telescoping the second part gives
$|\mathcal{C}_{K}\cap Q_{N}|=\sum_{m=1}^{M}\Bigl(((K-1)m+1)^{d}-((K-1)m)^{d}\Bigr)+(N-M)^{d}.$
Using $(x+1)^{d}-x^{d}=dx^{d-1}+O_{d}(x^{d-2})$ and $\sum_{m\leq M}m^{d-1}=\frac{M^{d}}{d}+O_{d}(M^{d-1})$, we obtain
$\sum_{m=1}^{M}\Bigl(((K-1)m+1)^{d}-((K-1)m)^{d}\Bigr)=(K-1)^{d-1}M^{d}+O_{d,K}(M^{d-1}).$
Therefore
$\frac{|\mathcal{C}_{K}\cap Q_{N}|}{N^{d}}=(K-1)^{d-1}\Bigl(\frac{M}{N}\Bigr)^{d}+\Bigl(1-\frac{M}{N}\Bigr)^{d}+O_{d,K}\Bigl(\frac{1}{N}\Bigr).$
Since $M/N\to 1/K$, the limit equals
$(K-1)^{d-1}\Bigl(\frac{1}{K}\Bigr)^{d}+\Bigl(1-\frac{1}{K}\Bigr)^{d}=\Bigl(1-\frac{1}{K}\Bigr)^{d-1}.\qed$
@@
Let $\delta:=\overline{d}(A)>0$. Choose $K$ so that
$1-\Bigl(1-\frac{1}{K}\Bigr)^{d-1}<\frac{\delta}{4}.$
By Lemma 2.12, there exists $N_{0}$ such that for all $N\geq N_{0}$,
$\frac{|Q_{N}\setminus\mathcal{C}_{K}|}{|Q_{N}|}<\frac{\delta}{2}.$
Pick $N_{i}\to\infty$ with $|A\cap Q_{N_{i}}|/|Q_{N_{i}}|\to\delta$ and $N_{i}\geq N_{0}$, so for large $i$, $|A\cap Q_{N_{i}}|/|Q_{N_{i}}|>3\delta/4$. Then
$\frac{|A\cap\mathcal{C}_{K}\cap Q_{N_{i}}|}{|Q_{N_{i}}|}\geq\frac{|A\cap Q_{N_{i}}|}{|Q_{N_{i}}|}-\frac{|Q_{N_{i}}\setminus\mathcal{C}_{K}|}{|Q_{N_{i}}|}\geq\frac{3\delta}{4}-\frac{\delta}{2}=\frac{\delta}{4}.$
Taking $\limsup_{i}$ gives $\overline{d}(A\cap\mathcal{C}_{K})\geq\delta/4>0$. ∎